A line is such that  its segment between the straight lines 5x-y-4=0 and 3x-4y-4=0 is bisected at the point (1,5).obtain the equation

Grade:12

1 Answers

Vikas TU
7 years ago
assume the eqn. of line be y = mx + c
it has two constants m and c .
after passing the points from (1,5) the eqn. becomes,
5 = m + c ….........................................(1)
let this line intersecting the given lines  5x-y-4=0 and 3x-4y-4=0
at (x1, 5x1 – 4)  and  (x2, (3x2-4)/4)
Now use the distance formulae from (x1, 5x1 – 4) to (1,5) and 
from (x2, (3x2-4)/4) to (1,5) and equate them.